OpenStudy (anonymous):

If C = {2, 4, 6, 8, 10, 12, 14, 16} and D = {numbers divisible by 4 from 1 to 20}, what is C ∩ D?

OpenStudy (anonymous):

from 1 to 20 numbers divisible by 4 are: 4 ,8 , 12 , 16, 20 \[\LARGE \frac44=1 , \quad \frac{8}{4}=2 ,\quad \frac{12}{4}=3 ,\quad \frac{16}{4}=4 ,\quad \frac{20}{4}=5\] so \[\LARGE C = \{2, 4, 6, 8, 10, 12, 14, 16\}\] and \[\LARGE D=\{1,2,3,4,5\}\] \[\LARGE C\cap D=\{2,4,\}\] ;)

OpenStudy (anonymous):

@kcute_face_bhunnie-_- Do you understand it ?

OpenStudy (anonymous):

D={4,8,12,16,20} \[ C\cap D=\{ 4,8,12,16 \} \]